(10pts) 2. Solve for x: log6(x + 5) + logox = 2 (10pts) 3. Solve for x: log(7x) +log(x - 2) = log5
Answer #1

Date: } for log ents) + log ned legat logbo legles) да с мае в“ (21+5) = 2 ) = 22 (0+5) = 6 n (22+5)= 36 A ==4 log (170) + lo